OpenStudy (anonymous):

Evaluate and write your answer in scientific notation: (2*10^2)(1.3*10^-3) @Calculus1

OpenStudy (anonymous):

\[2.6*10^{-1} \]

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

well whats this >>>Evaluate the following. Write your answer in scientific notation. (3*10^7)(6.1*10^3)

OpenStudy (anonymous):

\[(2*10{}^{\wedge}2)(1.3*10{}^{\wedge}-3) \]\[(200)\left(1.3*\frac{1}{1000}\right) =\]\[260*\frac{1}{1000}=\frac{260}{1000}=0.26 = 2.6*\frac{1}{10}=2.6 * 10^{-1} \]

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

Thank you for the medal.